OpenStudy (anonymous):

Line A is perpendicular to the line with equation y = -1/3 x and the point is on line A . Which of the following is the equation of line A

OpenStudy (anonymous):

Can you tell me the slope of this line y=(-1/3)x? hint: y=mx+c is the general equation of straight line with slope m and y-intercept c

OpenStudy (anonymous):

When we we are given an equation of a line and asked to find a line perpendicular to it what we should do is to take the slope in your case is -1/3x and get its negative reciprocal which is -(1/m) since your m=-1/3 its negative reciprocal is m=3 that negative reciprocal which is 3 is now you slope for the line you are trying to find the equation of line A is then y=3x+b b being the y intercept of that line

OpenStudy (anonymous):

@jacjie

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."